OpenStudy (anonymous):

EASY QUESTION! WILL GIVE MEDAL AND BECOME A FAN! PLEASE HELP?! Which of the following are measurements of the sides of a right triangle? Select one: a. 3, 4, 5 b. 12, 26, 28 c. 6, 14, 17 d. none of these

OpenStudy (mary.rojas):

... A based on A^2+B^2=C^2 A=3 B=4 C=5 3^2+4^2=5^2..... 9+16=25 SO A is correct

OpenStudy (nurali):

Use the theorem of Pythagoras: the square of the long side should be equal to the sum of the squares of the short sides. 5^2 = 3^2 + 4^2